### v1

@@ Appendix A Proof of Results in Section 2
A.1 Proof of Lemma 2.1
{Proof}
To prove Lemma 2.1, we introduce the kernel defined as $A(K^{\prime},K):=P^{K^{\prime}}_{K}-P_{K}$, where $P^{K^{\prime}}_{K}:=Q_{K^{\prime}}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})$ with $K^{\prime},K\in\mathcal{S}$, which works as a bridge to connect $P_{K}$ and $P_{K^{\prime}}$. Then we have:
(46) $\begin{split}A(K^{\prime},K)&=Q+K^{\prime\top}RK^{\prime}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})-P_{K}\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~{}~{}~{}~{}+(K^{\prime}-K)^{\top}\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)\\ &~{}~{}~{}~{}+\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)^{\top}(K^{\prime}-K)\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &=+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K).\end{split}$
@@
(47) $\begin{split}A(K^{\prime},K)&=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~{}~{}~{}~{}+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K)\\ &=(\underbrace{K^{\prime}-K+(R+B^{\top}P_{K}B)^{-1}E_{K}}_{=:\mathcal{A}})^{\top}(R+B^{\top}P_{K}B)\mathcal{A}\\ &~{}~{}~{}~{}-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}\\ &\succeq-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K},\end{split}$
@@
(48) $P_{K}-P_{K^{\prime}}=\sum_{t=0}^{\infty}(A_{K}^{t})^{\top}A(K,K^{\prime})A_{K}^{t}.$
@@
(49) $\begin{split}C(K)-C(K^{*})&=\mathrm{Tr}\left((P_{K}-P^{*})\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(-\left(\sum_{t=0}^{\infty}(A_{K^{*}}^{t})^{\top}A(K^{*},K)A_{K^{*}}^{t}\right)\Sigma_{w}\right)\\ &\leq\mathrm{Tr}\left(\left(\sum_{t=0}^{\infty}(A_{K^{*}}^{t})^{\top}E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}A_{K^{*}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}\Sigma_{K^{*}}\right).\\ \end{split}$
@@ A.2 Proof of Lemma 12
(50) $\begin{split}&~{}~{}~{}~{}~{}C(K^{\prime})-C(K)\\ &=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left(\sum_{t=0}^{\infty}(A_{K^{\prime}}^{t})^{\top}A({K^{\prime}},K)A_{K^{\prime}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left[({K^{\prime}}-{K})^{\top}(R+B^{\top}P_{K}B)({K^{\prime}}-{K})+({K^{\prime}}-{K})^{\top}E_{K}+E_{K}^{\top}({K^{\prime}}-{K})\right]\Sigma_{K^{\prime}}\right)\end{split}$
@@
A.3 Proof of Lemma 2.2
{Proof}
@@
(51) $\lVert P_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})},~{}~{}~{}\lVert\Sigma_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(Q)}.$
@@
From [10, Lemma 17], we can derive the upper bound of $\lVert\mathcal{T}_{K}\rVert$ as
@@
Define the identity operator $\mathcal{I}(X):=X$ Then we have the following results from [10, Lemma 18]:
@@
Then from [10, Lemma 16], we can prove the result of perturbation of $\Sigma_{K}$, as stated in Lemma 2.2
A.4 Proof of Lemma 2.2
{Proof}
When $\lVert K-K^{\prime}\rVert=\min\{h(C(K)),\lVert K\rVert\}$, by [10, Lemma 24], an upper bound on $\lVert P_{K^{\prime}}-P_{K}\rVert$ can be derived as:
(52) $\begin{split}\lVert P_{K^{\prime}}-P_{K}\rVert\leq{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2\lVert K\rVert^{2}\lVert R\rVert\lVert B\rVert+\lVert K\rVert\lVert R\rVert\right)}\lVert K-K^{\prime}\rVert\end{split}$
@@
(53) $\begin{split}C(K)-C(K^{*})&\geq C(K)-C(K^{\prime})\\ &=\mathrm{Tr}\left((P_{K}-P_{K^{\prime}})\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(-\left(\sum_{t=0}^{\infty}(A_{K^{\prime}}^{t})^{\top}A(K^{\prime},K)A_{K^{\prime}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left(\sum_{t=0}^{\infty}(A_{K^{\prime}}^{t})^{\top}E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}A_{K^{\prime}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}P_{K^{\prime}}^{w}\right)\\ &\geq\frac{\sigma_{1}(\Sigma_{w})}{\lVert R+B^{\top}P_{K}B\rVert}\mathrm{Tr}\left(E_{K}^{\top}E_{K}\right)\\ \end{split}$
@@
(54) $\begin{split}\lVert K\rVert&\leq\frac{1}{\lambda_{1}(R)}\left(\sqrt{\frac{(C(K)-C(K^{*}))\lVert R+B^{\top}P_{K}B\rVert}{\lambda_{1}(\Sigma_{w})}}+\lVert B^{\top}P_{K}A\rVert\right)\\ &\leq\underbrace{\frac{1}{\lambda_{1}(R)}\left(\sqrt{\frac{(C(K)-C(K^{*}))\left(\lVert R\rVert+\lVert B\rVert^{2}\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\right)}{\lambda_{1}(\Sigma_{w})}}+\lVert B\rVert\lVert A\rVert\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\right)}_{=:b_{K}(C(K))}\\ \end{split}$
@@
(55) $\begin{split}&~{}~{}~{}|C(K^{\prime})-C(K)|=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\leq\lVert P_{K^{\prime}}-P_{K}\rVert\mathrm{Tr}(\Sigma_{w})\\ &\leq\underbrace{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2b^{2}_{K}(C(K))\lVert R\rVert\lVert B\rVert+b_{K}(C(K))\lVert R\rVert\right)\mathrm{Tr}(\Sigma_{w})}_{=:h_{C}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$
A.5 Proof of Lemma 2.2
{Proof}
@@
(56) $\begin{split}\lVert\nabla C(K)\rVert&\leq\sqrt{4\lVert\Sigma_{K}\rVert^{2}\mathrm{Tr}\left(E_{K}^{\top}E_{K}\right)}\\ &\leq\underbrace{\sqrt{4\left(\frac{C(K)}{\lambda_{1}(Q)}\right)^{2}\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}}_{=:b_{\nabla}(C(K))}.\end{split}$
Under the condition on $K^{\prime}$: $\lVert K-K^{\prime}\rVert\leq h(C(K))$, then the second term is bounded by:
(57) $\begin{split}&~{}~{}~{}~{}\lVert 2E_{K}(\Sigma_{K}-\Sigma_{K^{\prime}})\rVert\\ &\leq\underbrace{2\sqrt{\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}h_{\Sigma}{(C(K))}}_{=:\alpha_{1}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$
@@
(58) $\begin{split}&\lVert 2(E_{K^{\prime}}-E_{K})\Sigma_{K^{\prime}}\rVert\leq 2\lVert E_{K^{\prime}}-E_{K}\rVert\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\\ &\leq 2\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{0})}}\right)\lVert K-K^{\prime}\rVert\\ &~{}~{}~{}+2\alpha_{2}(C(K))\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert B\rVert\lVert A\rVert+\lVert B\rVert^{2}\lVert K\rVert\right)\lVert K-K^{\prime}\rVert\\ &\leq\alpha_{3}(C(K))\lVert K-K^{\prime}\rVert\end{split}$
@@
(59) $\lVert\nabla C(K^{\prime})-\nabla C(K)\rVert\leq\underbrace{(\alpha_{1}(C(K))+\alpha_{3}(C(K)))}_{=:h_{\nabla}(C(K))}\lVert K-K^{\prime}\rVert$
@@ Appendix B Proof of Results in Section 3
B.1 Proof of Theorem 3.1
{Proof}
Using Lemma 12 and following the step by [10, Lemma 21], if
(60) $\begin{split}\eta_{i}\leq\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{\lVert B\rVert\lVert\nabla C(K_{i})\rVert(1+\lVert A+BK_{i}\rVert)},\right.\\ \left.\frac{\lambda_{1}(Q)}{2C(K_{i})\lVert R+B^{\top}P_{K_{i}}B\rVert}\right\},\end{split}$
@@
(61) $C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$
@@
(62) $\eta_{i}\leq\underbrace{\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{2\lVert B\rVert b_{\nabla}(C(K_{i}))},\frac{\lambda_{1}(Q)}{2C(K_{i})\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}{(\Sigma_{w})}}\right)}\right\}}_{=:h_{\mathrm{GD}}(C(K))},$
where $b_{\nabla}$ is defined in (56). From (61), we know that $C(K_{i+1})\leq C(K_{i})$. Then we have $h_{\mathrm{GD}}(C(K_{i+1}))\geq h_{\mathrm{GD}}(C(K_{i}))$. The remainder of the proof proceeds by setting $\eta_{i}=\eta_{0},\forall i\in\mathbb{Z}_{+}$, which represents the lower bound of the convergence rate, and then following the steps outlined in [10, Lemma 22].
@@ B.2 Update of Natural Policy Gradient
(63) $u_{t}=\underbrace{\left[\begin{array}[]{cccc}x_{t}^{\top}&0&...&0\\ 0&x_{t}^{\top}&...&...\\ ...&...&...&0\\ 0,&...&0&x_{t}^{\top}\end{array}\right]}_{=:X_{t}\in\mathbb{R}^{n_{u}\times n_{x}n_{u}}}\left[\begin{array}[]{cc}k_{1}^{\top}\\ k_{2}^{\top}\\ ...\\ k_{n_{x}}^{\top}\end{array}\right]=X_{t}\bar{K}=Kx_{t}$

### v2

@@ Appendix A Proof of Results in Section 2
A.1 Proof of Lemma 1
Proof 1
To prove Lemma 1, we introduce the kernel defined as $A(K^{\prime},K):=P^{K^{\prime}}_{K}-P_{K}$, where $P^{K^{\prime}}_{K}:=Q_{K^{\prime}}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})$ with $K^{\prime},K\in\mathcal{S}$, which works as a bridge to connect $P_{K}$ and $P_{K^{\prime}}$. Then we have:
$\begin{split}A(K^{\prime},K)&=Q+K^{\prime\top}RK^{\prime}+(A+BK^{\prime})^{\top}P_{K}(A+BK^{\prime})-P_{K}\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~~~~+(K^{\prime}-K)^{\top}\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)\\ &~~~~+\left((R+B^{\top}P_{K}B)K+B^{\top}P_{K}A\right)^{\top}(K^{\prime}-K)\\ &=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &=+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K).\end{split}$ (46)
@@ A.1 Proof of Lemma 1
$\begin{split}A(K^{\prime},K)&=(K^{\prime}-K)^{\top}(R+B^{\top}P_{K}B)(K^{\prime}-K)\\ &~~~~+(K^{\prime}-K)^{\top}E_{K}+E_{K}^{\top}(K^{\prime}-K)\\ &=(\underbrace{K^{\prime}-K+(R+B^{\top}P_{K}B)^{-1}E_{K}}_{=:\mathcal{A}})^{\top}(R+B^{\top}P_{K}B)\mathcal{A}\\ &~~~~-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K}\\ &\succeq-E_{K}^{\top}(R+B^{\top}P_{K}B)^{-1}E_{K},\end{split}$ (47)
@@
$P_{K}-P_{K^{\prime}}=\sum_{t=0}^{\infty}(A_{K}^{t})^{\top}A(K,K^{\prime})A_{K}^{t}.$ (48)
@@
[…]
@@ A.2 Proof of Lemma 2
$\begin{split}&~~~~~C(K^{\prime})-C(K)\\ &=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left(\sum_{t=0}^{\infty}(A_{K^{\prime}}^{t})^{\top}A({K^{\prime}},K)A_{K^{\prime}}^{t}\right)\Sigma_{w}\right)\\ &=\mathrm{Tr}\left(\left[({K^{\prime}}-{K})^{\top}(R+B^{\top}P_{K}B)({K^{\prime}}-{K})+({K^{\prime}}-{K})^{\top}E_{K}+E_{K}^{\top}({K^{\prime}}-{K})\right]\Sigma_{K^{\prime}}\right)\end{split}$ (50)
@@
A.3 Proof of Lemma 3
Proof 3
@@ A.3 Proof of Lemma 3
$\lVert P_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})},~~~\lVert\Sigma_{K}\rVert\leq\frac{C(K)}{\lambda_{1}(Q)}.$ (51)
@@
From [9, Lemma 17], we can derive the upper bound of $\lVert\mathcal{T}_{K}\rVert$ as
@@
Define the identity operator $\mathcal{I}(X):=X$ Then we have the following results from [9, Lemma 18]:
@@
Then from [9, Lemma 16], we can prove the result of perturbation of $\Sigma_{K}$, as stated in Lemma 3
A.4 Proof of Lemma 4
Proof 4
When $\lVert K-K^{\prime}\rVert=\min\{h(C(K)),\lVert K\rVert\}$, by [9, Lemma 24], an upper bound on $\lVert P_{K^{\prime}}-P_{K}\rVert$ can be derived as:
$\begin{split}\lVert P_{K^{\prime}}-P_{K}\rVert\leq{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2\lVert K\rVert^{2}\lVert R\rVert\lVert B\rVert+\lVert K\rVert\lVert R\rVert\right)}\lVert K-K^{\prime}\rVert\end{split}$ (52)
@@ A.4 Proof of Lemma 4
[…]
@@
[…]
@@
$\begin{split}&~~~|C(K^{\prime})-C(K)|=\mathrm{Tr}\left((P_{K^{\prime}}-P_{K})\Sigma_{w}\right)\leq\lVert P_{K^{\prime}}-P_{K}\rVert\mathrm{Tr}(\Sigma_{w})\\ &\leq\underbrace{6\left(\frac{C(K)}{\lambda_{1}(\Sigma_{w})\lambda_{1}(Q)}\right)^{2}\left(2b^{2}_{K}(C(K))\lVert R\rVert\lVert B\rVert+b_{K}(C(K))\lVert R\rVert\right)\mathrm{Tr}(\Sigma_{w})}_{=:h_{C}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$ (55)
A.5 Proof of Lemma 5
Proof 5
@@ A.5 Proof of Lemma 5
$\begin{split}\lVert\nabla C(K)\rVert&\leq\sqrt{4\lVert\Sigma_{K}\rVert^{2}\mathrm{Tr}\left(E_{K}^{\top}E_{K}\right)}\\ &\leq\underbrace{\sqrt{4\left(\frac{C(K)}{\lambda_{1}(Q)}\right)^{2}\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}}_{=:b_{\nabla}(C(K))}.\end{split}$ (56)
[…]
$\begin{split}&~~~~\lVert 2E_{K}(\Sigma_{K}-\Sigma_{K^{\prime}})\rVert\\ &\leq\underbrace{2\sqrt{\frac{(C(K)-C(K^{*}))}{\sigma_{1}(\Sigma_{w})}\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{w})}}\right)}h_{\Sigma}{(C(K))}}_{=:\alpha_{1}(C(K))}\lVert K-K^{\prime}\rVert.\end{split}$ (57)
@@
$\begin{split}&\lVert 2(E_{K^{\prime}}-E_{K})\Sigma_{K^{\prime}}\rVert\leq 2\lVert E_{K^{\prime}}-E_{K}\rVert\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\\ &\leq 2\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K)}{\lambda_{1}{(\Sigma_{0})}}\right)\lVert K-K^{\prime}\rVert\\ &~~~+2\alpha_{2}(C(K))\left(\lVert\Sigma_{K}\rVert+\frac{C(K)}{\lambda_{1}(Q)}\right)\left(\lVert B\rVert\lVert A\rVert+\lVert B\rVert^{2}\lVert K\rVert\right)\lVert K-K^{\prime}\rVert\\ &\leq\alpha_{3}(C(K))\lVert K-K^{\prime}\rVert\end{split}$ (58)
@@
$\lVert\nabla C(K^{\prime})-\nabla C(K)\rVert\leq\underbrace{(\alpha_{1}(C(K))+\alpha_{3}(C(K)))}_{=:h_{\nabla}(C(K))}\lVert K-K^{\prime}\rVert$ (59)
@@ Appendix B Proof of Results in Section 3
B.1 Proof of Theorem 1
Proof 6
Using Lemma 2 and following the step by [9, Lemma 21], if
$\begin{split}\eta_{i}\leq\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{\lVert B\rVert\lVert\nabla C(K_{i})\rVert(1+\lVert A+BK_{i}\rVert)},\right.\\ \left.\frac{\lambda_{1}(Q)}{2C(K_{i})\lVert R+B^{\top}P_{K_{i}}B\rVert}\right\},\end{split}$ (60)
@@ B.1 Proof of Theorem 1
$C(K_{i+1})-C(K^{*})\leq\left(1-\frac{2\eta_{i}\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{\lVert\Sigma_{K}^{*}\rVert}\right)(C(K_{i})-C(K^{*})).$ (61)
@@
$\eta_{i}\leq\underbrace{\frac{1}{32}\min\left\{\left(\frac{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}{C(K_{i})}\right)^{2}\frac{1}{2\lVert B\rVert b_{\nabla}(C(K_{i}))},\frac{\lambda_{1}(Q)}{2C(K_{i})\left(\lVert R\rVert+\frac{\lVert B\rVert^{2}C(K_{i})}{\lambda_{1}{(\Sigma_{w})}}\right)}\right\}}_{=:h_{\mathrm{PGD}}(C(K))},$ (62)
where $b_{\nabla}$ is defined in (56). From (61), we know that $C(K_{i+1})\leq C(K_{i})$. Then we have $h_{\mathrm{PGD}}(C(K_{i+1}))\geq h_{\mathrm{PGD}}(C(K_{i}))$. The remainder of the proof proceeds by setting $\eta_{i}=\eta_{0},\forall i\in\mathbb{Z}_{+}$, which represents the lower bound of the convergence rate, and then following the steps outlined in [9, Lemma 22].
@@ B.2 Update of Natural Policy Gradient
$u_{t}=\underbrace{\left[\begin{array}[]{cccc}x_{t}^{\top}&0&...&0\\ 0&x_{t}^{\top}&...&...\\ ...&...&...&0\\ 0,&...&0&x_{t}^{\top}\end{array}\right]}_{=:X_{t}\in\mathbb{R}^{n_{u}\times n_{x}n_{u}}}\left[\begin{array}[]{cc}k_{1}^{\top}\\ k_{2}^{\top}\\ ...\\ k_{n_{x}}^{\top}\end{array}\right]=X_{t}\bar{K}=Kx_{t}$ (63)